OpenStudy (anonymous):

Solve this system of equations using the addition method. 6a - 5b =9 2a + 5b = 23 @Bea_26

OpenStudy (anonymous):

If you notice, you can see that the 5b's can be eliminated because they are opposite signs.. When you eliminate the 5bs, what do you have left?

OpenStudy (anonymous):

Make sure when you see what you have left, you are adding the rest of the equation to each other. That's because the only way we eliminated the 5b was through addition, so don't forget that!!

OpenStudy (anonymous):

i dont get u

OpenStudy (anonymous):

Okay, I tried.. :(

OpenStudy (unklerhaukus):

6a - 5b =9 2a + 5b = 23 add the equations like this ( 6a - 5b ) + ( 2a + 5b )= ( 9 )+( 23 ) 6a + 2a - 5b + 5b = 9 + 23

OpenStudy (anonymous):

then wht do i do

OpenStudy (unklerhaukus):

simplify

OpenStudy (anonymous):

do i simplify the whole thing

OpenStudy (unklerhaukus):

you are solving for a this could be a first step (6 + 2)a + (- 5 + 5)b = 9 + 23

OpenStudy (texaschic101):

whats 6a + 2a ? whats - 5b + 5b ? whats 9 + 23 ? add these and what do you get ?

OpenStudy (anonymous):

6a+2a=8a 5b+5b=10b 9+23=32 and then what do i do

OpenStudy (unklerhaukus):

* -5b+5b=

OpenStudy (anonymous):

wht was tht @UnkleRhaukus

OpenStudy (texaschic101):

check this one again...-5a + 5b.....they cancel each other out leaving 0 let me rephrase this.... 6a - 5b = 9 2a + 5b = 23 ------------ 8a + 0 = 32 8a = 32

OpenStudy (unklerhaukus):

one of the five b's is negative

OpenStudy (anonymous):

@texaschic101 so whts the answer

OpenStudy (anonymous):

You solve them simultaneously by multiplying (2a+5b=23) by -3 so that the coefficient of a becomes -6 and then a is eliminated; (2a+5b=23)*-3 >-6a-15b=-69...(1) 6a-5b=9 -6a-15b=-69 -20b=-60 b=3 6a-5(3)=9 a=-1 Am I Correct.?

OpenStudy (texaschic101):

right now, we have this... 8a =32 you want to get x by itself, so now you divide both sides by 8 can you do that ?

OpenStudy (texaschic101):

@Bea_26 ....she has to do addition property, not multiplication

OpenStudy (texaschic101):

did you figure it out @shorty17 ?

OpenStudy (anonymous):

@texaschic101 it addition

OpenStudy (anonymous):

Well sorry smarty pants Texas.. whatever your name is.

OpenStudy (anonymous):